OpenStudy (anonymous):

(3x^-2 y^3)/ (9x^3 y^-5) =1)(3x^-5)/y^8 2)(xy^2)/3 3) 3xy^2 4)(y^6)/ 3x^5

OpenStudy (anonymous):

x^n/ x^m = x^n-m

OpenStudy (anonymous):

??

OpenStudy (anonymous):

That's the rule to use. Just simplify.

OpenStudy (anonymous):

12x^1 that wat i got

OpenStudy (anonymous):

u just confuzed me more

OpenStudy (anonymous):

I'm getting y^8/(3x^5) 3/9 reduced to 1/3 x^-2/x^3 reduces to x^-5 or 1/x^5 y^3/y^-5 reduces to y^8

OpenStudy (anonymous):

You should get y^8/3x

OpenStudy (anonymous):

sorry 3x^5 in the denom

OpenStudy (anonymous):

the answer clostest to that has 3x^5 not 3x

OpenStudy (dumbcow):

are you sure the question is correct... i don't agree with any of the answers \[\frac{3x^{-2}y^{3}}{9x^{3}y^{-5}} = \frac{y^{3+5}}{3x^{3+2}} = \frac{y^{8}}{3x^{5}}\]

OpenStudy (anonymous):

Okay. 3*x^-2*y^3 3 x^-2 y^3 ____________=__ * _____ * ______ 9*x^3*y^-5 9 x^3 y^-5 Reduce each of those 3 fractions using the rule I gave to get 1/3 * x^-5 * y^8

OpenStudy (anonymous):

Gotta go. Hope that explains it.